OpenStudy (anonymous):

Solve . sin(^2)x-3sinxcosx+2cos(^2)x=0

OpenStudy (anonymous):

try factoring

OpenStudy (anonymous):

(Sin(x) - Cos(x) ) (Sin(x) - 2 Cos(x))=0

OpenStudy (anonymous):

Sin(x) - Cos(x) =0 Sin(x)= Cos(x) that's true when x is Pi/4 + Pi n (n=1,2,3,4)

OpenStudy (anonymous):

(Sin(x) - 2 Cos(x))=0 Sin(x)=2 Cos(x) tan(x)=2 arctan(2)

OpenStudy (anonymous):

Thanks!